OpenStudy (juan1857):

Express 5√3-5i in polar form.

OpenStudy (mww):

this expression is a bit ambiguous. Need to write it a bit better. Try the equation tool.

OpenStudy (juan1857):

ok so \[5\sqrt{3} -5i\]

OpenStudy (amorfide):

\[z=a+bi\] \[r=|z|=\sqrt{a^{2}+b^{2}}\] \[\theta=\tan^{-1}(\frac{ b }{ a })\] polar coordinates are in the form of \[(r,\theta)\]

OpenStudy (amorfide):

so let \[z=5\sqrt{3}-5i\] a is the real part, y is the imaginary part so you have \[r=|z|=\sqrt{(5\sqrt{3})^{2}+(-5)^{2}}\]

OpenStudy (amorfide):

you can finish the rest using this

OpenStudy (amorfide):

I meant b is the imaginary part, not y, sorry

OpenStudy (juan1857):

for some reason i got 40 for r, I don't know if its correct

OpenStudy (amorfide):

you should get 25(3)+25=100 then sqrt 100= 10

OpenStudy (amorfide):

if you write out your calculation step by step I can see where you went wrong

OpenStudy (juan1857):

do I multiply the square root with 5 ?

OpenStudy (amorfide):

yes

OpenStudy (amorfide):

it is 5 lots of sqrt3 so 5sqrt3 you square it to get 5sqrt3 x 5sqrt3 which gives (5x5)(sqrt3 x sqrt3) which gives 25sqrt9 sqrt9=3 25(3)=75

OpenStudy (juan1857):

ah, now I see, Thank you Very much

OpenStudy (amorfide):

Continue through the question to make sure you get the correct final answer

OpenStudy (juan1857):

for theta I got -60, but I think that wouldn't work.

OpenStudy (amorfide):

you shouldn't get 60 at all

OpenStudy (amorfide):

\[\theta = \tan^{1}(\frac{ b }{ a })= \tan^{-1}(\frac{ 5 }{ 5\sqrt{3} })\] 

OpenStudy (amorfide):

you should get 30 degrees

OpenStudy (amorfide):

1/6pi

OpenStudy (juan1857):

sorry, I am still getting 60, I know your write because I know the answer but I don't know how

OpenStudy (amorfide):

show me your working out

OpenStudy (amorfide):

you are doing \[\tan^{-1}(\frac{ 5\sqrt{3} }{ 5 })\]

OpenStudy (amorfide):

switch the fraction around

OpenStudy (amorfide):

and you get 30

OpenStudy (juan1857):

I am doing that,and still get 60. Maybe its the calculator

OpenStudy (amorfide):

do \[\frac{ 5 }{ 5\sqrt{3} }\]

OpenStudy (amorfide):

\[\tan^{-1}(ans)\]

OpenStudy (juan1857):

man I'm still getting the answer. I don't want to take more of your time. Thanks!

OpenStudy (amorfide):

wait maybe try \[5\sqrt{3}\] press equals then \[\frac{ 5 }{ ans }\] press equals then \[\tan^{-1}(ans)\] if this doesn't work then I give up xD

OpenStudy (juan1857):

oh so now I got it thanks man!